How did the tempruature change?
First it increased by 5%, and then it increased by 20%

Answer by ikleyn(52781)

.
How did the tempruature change?
First it increased by 5%, and then it increased by 20%
~~~~~~~~~~~~~~~~~~~~

Let T be the initial temperature.


Then after first increase the temperature becomes  T*(1+0.05) = 1.05T.


After the second increase, the temperature becomes 1.05T*(1+0.2)T = 1.05*1.2T = 1.26T.


The final increase is 26%, comparing to initial temperature.

Solved.

----------------------

If you want to see many other similar  (and different)  solved problems,  look into the lesson
    - Percentage problems on chains of discounts
in this site.